# Transferable Post-training via Inverse Value Learning

### v1

@@ 6 Discussions
Table 2: Comparison between Llama-2-70B based model guided by the standard 1.1B value model, the 1.1B model trained with a curriculum across 1.1B, 7B, and 13B models, and the model trained with a L1 normalization term.
Given that all models in the Llama-2 series share the same vocabulary, we designed a curriculum to train the model sequentially on the 1.1B, 7B, and 13B versions, and then evaluated its generalization performance on the powerful 70B model. We maintained consistent training data volume and iteration counts across experiments. As shown in Table 2, our findings indicate that this multi-model curriculum training can also mitigate over-fitting to some extent, functioning similarly to regularization techniques. However, this method necessitates longer training periods and more complex model scheduling processes. Consequently, it does not offer significant advantages in weak-to-strong generation over simpler normalization methods in our observation.
Figure 4: MT-Bench scores of Llama-3-8B fully fine-tuned, Llama-3-8B-base plugged with the TinyLlama 1.1B value model (cross vocabulary), and Llama-3-8B-base plugged with the Llama 3.1 value models (same vocabulary).
@@
We explored the transferability of model capabilities across different model families. We transferrs 1.1B parameter value models to Llama-3 using vocabulary mapping at decoding time, comparing it with Llama-3.2-1B guided and full fine-tuning. Interestingly, among all settings of the trained value model, only the curriculum-trained value model yielded acceptable generation ability in the cross-model and cross-vocabulary setting. However, as Figure 4 shows, while the transferred model grasped basic conversational abilities, it still underperformed compared to same-vocabulary value models and full fine-tuning. In addition to the generation ability of the value model, we attribute this performance gap partly to tokenization differences: source tokens can be mapped to multiple consecutive target tokens, which are unfamiliar to the source model in the following decoding steps, potentially degrading performance during decoding. While we’ve demonstrated the feasibility of cross-vocabulary value transfer, improving vocabulary mapping algorithms and cross-family model guidance remain promising areas for future research.
@@ 7 Conclusions
In this paper, we introduced a novel framework for transferable post-training via inverse value learning, which models residual logits adjustments using a separate value network. We systematically explored various aspects of this paradigm, including network architectures, pre-training weights, and connection schemes. Additionally, we addressed challenges related to overfitting and cross-vocabulary transfer by incorporating regularization techniques and a vocabulary mapping algorithm. Our results confirm that the value network exhibits broad transferability across models of different sizes and families, highlighting the potential of logits-space adjustments for efficient model adaptation.
@@ A.1 Probe Architecture
In this work, we adapt the MLP layer from the LLaMA architecture as a linear probe, which can be formally expressed as:
$\mathbf{y}=\mathbf{W}_{\text{gate}}(\sigma(\mathbf{W}_{\text{gate}}\mathbf{x}+\mathbf{b}_{\text{gate}})\odot(\mathbf{W}_{\text{up}}\mathbf{x}+\mathbf{b}_{\text{up}}))+\mathbf{b}_{\text{gate}}$

### v2

@@ 6 Discussions
Table 2: Comparison between Llama-2-70B base model guided by the standard 1.1B value model, the 1.1B model trained with a curriculum across 1.1B, 7B, and 13B models, and the model trained with a L1 normalization term.
Given that all models in the Llama-2 series share the same vocabulary, we designed a curriculum to train an 1.1B value model sequentially on the 1.1B, 7B, and 13B versions, and then evaluated its generalization performance on the powerful 70B model. We maintained consistent training data volume and iteration counts across experiments. As shown in Table 2, our findings indicate that this multi-model curriculum training can also mitigate over-fitting to some extent, functioning similarly to regularization techniques. However, this method necessitates longer training periods and more complex model scheduling processes. Consequently, it does not offer significant advantages in weak-to-strong generation over simpler normalization methods in our observation.
Figure 4: MT-Bench scores of Llama-3-8B fully fine-tuned, Llama-3-8B-base plugged with the TinyLlama 1.1B value model (cross vocabulary), and Llama-3-8B-base plugged with the Llama 3.2 value models (same vocabulary).
@@
We further explored the transferability of model capabilities across different model families. We transfers Llama-2-vocabulary-based 1.1B parameter value models to Llama-3 using vocabulary mapping at decoding time, comparing it with Llama-3.2-1B guided and full fine-tuning performance. Interestingly, among all settings of the trained value model, only the curriculum-trained value model yielded acceptable generation ability in the cross-model and cross-vocabulary setting. However, as Figure 4 shows, while the transferred model grasped basic conversational abilities, it still underperformed compared to same-vocabulary value models and full fine-tuning. In addition to the generation ability of the value model, we attribute this performance gap partly to tokenization differences: source tokens can be mapped to multiple consecutive target tokens, which are unfamiliar to the source model in the following decoding steps, potentially degrading performance during decoding. While we’ve demonstrated the feasibility of cross-vocabulary value transfer, improving vocabulary mapping algorithms and cross-family model guidance remain promising areas for future research.
@@ 7 Conclusions
In this paper, we introduced a novel method for transferable post-training called inverse value learning, which models residual logits adjustments using a separate value network. We systematically explored various aspects of this paradigm, including network architectures, pre-training weights, and connection schemes. Additionally, we addressed challenges related to overfitting and cross-vocabulary transfer by incorporating regularization techniques and a vocabulary mapping algorithm. Our results confirm that the value network exhibits broad transferability across models of different sizes and families, highlighting the potential of logits-space adjustments for efficient model adaptation.
@@
Acknowledgments
We sincerely thank the anonymous ARR reviewers for their insightful comments and valuable suggestions. This work was supported by Beijing Natural Science Foundation (L243006), the Basic Research Program of ISCAS (Grant No. ISCAS-ZD-202401), the Youth Talent Program of ISCAS (Grant No. SYQ2022-3).
@@ A.1 Probe Architecture
In this work, we adapt the MLP layer from the Llama (Touvron et al., 2023) architecture as a linear probe, which can be formally expressed as:
$\mathbf{y}=\mathbf{W}_{\text{down}}(\sigma(\mathbf{W}_{\text{gate}}\mathbf{x}+\mathbf{b}_{\text{gate}})\odot(\mathbf{W}_{\text{up}}\mathbf{x}+\mathbf{b}_{\text{up}}))+\mathbf{b}_{\text{down}}$
@@ A.2 Failure modes
Model Config #GPU Time/Peak Memory
128 tokens 512 tokens 1024 tokens
1.1B 1 1.99s/2.09GB 6.99s/2.09GB 13.36s/2.10GB
7B 1 2.87s/12.65GB 10.98s/12.84GB 22.06s/13.09GB
1.1B + 7B 1 5.13s/14.79GB 20.83s/15.33GB 49.74s/15.91GB
13B 1 3.48s/24.51GB 12.98s/24.79GB 26.09s/25.18GB
1.1B + 13B 1 5.73s/26.66GB 26.10s/27.30GB 61.61s/28.14GB
70B 4 11.76s/32.41GB 20.36s/32.42GB 40.36s/32.44GB
1.1B + 70B 4 16.12s/33.08GB 33.07s/33.31GB 75.37s/33.78GB
Table 3: Time and Peak Memory per GPU for different model configurations and generate sequence lengths.
@@ Appendix C Training Details
Appendix D Additional Results
Comparison with Proxy Tuning
As a complement to our related work analysis, we conduct performance comparisons with proxy tuning on a 7B model, with results shown in Table 4. The empirical results demonstrate that our IVL-trained model achieves comparable performance to proxy tuning on the 7B model. Furthermore, IVL demonstrates higher inference efficiency as it does not require an additional 1B reference model during inference.
Llama-2-7B+ 1.1B 1.1B+1.1B
MT-Bench 4.62 4.65
ARC 42.7 40
Hellaswag 55 50.4
MMLU 43.2 44.4
TruthfulQA 31.1 33.4
Winogrande 66.2 66.7
Table 4: Performance comparison between Inverse Value Learning (7B+1.1B) and Proxy Tuning (7B+1.1B+1.1B).
Inference Costs
As shown in Table 3, we conduct experiments to measure inference time for generating sequences of different lengths using various combinations of base models and value models. All experiments were performed using bfloat16 precision and flash-attention v2 for acceleration on NVIDIA H800-80G GPUs. For models smaller than 13B parameters and their combinations, pipeline parallelism was not employed. We did not implement any additional optimizations for value model inference. The reported results are averaged over 5 independent runs.